OpenStudy (anonymous):

Daniel wants to have a 90 average in his math class at the end of the year. He is trying to determine what he needs to get on his final exam, which accounts for 10% of his grade, for this to work. Tests are weighted 50% of the grade, and he currently has a 85 for his test average; quizzes are weighted 15% of his grade, and he currently has a 95 quiz average; homework is weighted 15% of his grade and he currently has a 98 homework average; and projects are weighted 10% of his grade and he currently has a 92 project average. What does he need to get on his final exam to get a 90% in the class

OpenStudy (kirbykirby):

85*0.5+95*0.15+98*0.15+92*0.1 = X is taking the percentage weight of each corresponding grade. This will give a mark out of 90, so X/90 Thus, to get 90 in the class, you need a total score 90/100 So, you determine X/90 + ?/10 = 90/100 which is a sample task of finding the missing value ?

OpenStudy (kirbykirby):

you should find that he requires 93.5% on the final, or 9.35/10

OpenStudy (anonymous):

so at the moment he has an 80.65% in the class then you calculate by adding 81(approximately) over 90

OpenStudy (kirbykirby):

It's more precise if you keep the decimals 80.65/90 (unless your teacher requested to round that).

OpenStudy (kirbykirby):

Because effectively if you round up at this stage, you are saying he'll only need 9/10 on the final. And if you do the weighted percentage calculation by considering the 9/10 grade, it won't be at 90%

OpenStudy (kirbykirby):

(You'll find that he'll get 89.65%, and not 90%... which in real life situations might not get rounded up to 90%.. )

OpenStudy (anonymous):

Just in case anyone needs this for future reference, your just basically calculating everything together and solving. Here's how I solved it; .10x + .50(85) + .15(95) + .15(98) + .10(92) = 90 0.1x + 42.5 + 14.25 + 14.7 + 9.2 = 90 0.1x + 80.65 = 90 -80.65 = -80.65 0.1x = 9.35 0.1 = 0.1 x = 93.5 So your answer is 93.5 from plugging in the values into the equation and solving. You got this guys!